I need to set up the equation for a problem like this The enrollemnt in trig class is eight less than enrollemnt in algebra class. There are 32 in both classes combined find the number
Answer by kunal_122(19)

x-8+x=32
2x-8=32
2x=40
x=20
enrolment in algebra is 20 and in trig is 20-8=12
and these 2 add to 32